# LLMs hallucinate graphs too: a structural perspective

### v2

@@ Abstract
It is known that LLMs do hallucinate, that is, they return incorrect information as facts. In this paper, we introduce the possibility to study these hallucinations under a structured form: graphs. Hallucinations in this context are incorrect outputs when prompted for well known graphs from the literature (e.g. Karate club, Les Misérables, graph atlas). These hallucinated graphs have the advantage of being much richer than the factual accuracy –or not– of a statement; this paper thus argues that such rich hallucinations can be used to characterize the outputs of LLMs. Our first contribution observes the diversity of topological hallucinations from major modern LLMs. Our second contribution is the proposal of a metric for the amplitude of such hallucinations: the Graph Atlas Distance, that is the average graph edit distance from several graphs in the graph atlas set. We compare this metric to the Hallucination Leaderboard, a hallucination rank that leverages $10,000$ times more prompts to obtain its ranking.
@@
Large language models, hallucination, graphs, edit distance.
@@ 1 Introduction
Large Language Models (LLMs) recently attracted a lot of attention, thanks to sustained research efforts and a large spectrum of envisioned applications. This triggers a growing demand for tools to test and analyze these expensive to set up and complex objects. In particular, methods to efficiently identify, differentiate, watermark LLMs, and methods to compare their accuracy and notably the potential presence of hallucinations [15, 20, 19] are devised. The common denominator of these methods is the will to efficiently extract information from the LLM under scrutiny.
@@
To achieve this information extraction, one can distinguish two broad categories of approaches: white box approaches that rely on exploiting a privileged access to the model internals (e.g. probits, activation patterns, source code, model hyperparameters and weights), and black box approaches that merely allow an auditor to interact with that model. White box approaches provide precise answers, but are not always available, for instance to an external auditor assessing a closed source LLM. In such cases, black box approaches are necessary, and constitute the focus of this paper.
@@
To conduct a black box audit of a LLM, a common approach is to rely on multiple choice questions (MCQ) to test the accuracy of its answers. Using MCQ circumvents the difficulty and subjectivity of evaluating (potentially large numbers of) free text answers. To that end, several MCQ datasets are readily available, and cover a wide range of knowledge and reasoning domains. For instance, MedMCQA [17] contains 182k questions of 4 choices assessing the target model general medical knowledge. Using this dataset yields 2 bits of information about the target model per sent request. This might reveal inefficient if the auditor has to pay for each request or if the auditor has to limit the rate at which requests are sent.
@@
In this paper, we explore a different approach using prompts willing to obtain graphs, rather than text. As graphs are structured, requesting graphs thus circumvents the problem of analyzing free form text answers. And most importantly, a single request to obtain a $n$ node graph from a LLM yields $O(n^{2})$ structured information bits from that model, enabling a much more efficient information extraction rate per request for the auditor. For instance, requesting the standard "Les Misérables" graph is much more efficient than issuing requests like "do Jean Valljean and Cosette both appear in a chapter?", for any two of the 77 characters of the book. Moreover, throughout the years, network science has produced a number of standard benchmark graphs already ingested by most LLMs, providing a readily available ground truth. This ingestion from LLMs comes from the fact that they are available as data online (e.g. in the SNAP [14] or KONECT [13] repositories among others), and as the process of training these LLMs is known to consume most of available data online [10].
While querying LLMs for graph structures is unexplored and might sound promising, it raises a number of questions we further investigate in the remainder of this paper. First, we detail the envisioned interaction framework in which the objective is to collect a graph as a LLM output. Once this basic interaction framework has been established, we first focus on the quality of the collected graphs. Namely, are output graphs accurate with regard to the original queried graph? And moreover, are some answers wrong and inaccurate, in other words do LLMs hallucinate graphs? Second, we showcase the potential of graph requests and their capacity to extract information regarding target LLMs in a request-efficient way. To that end, we compare our proposal (the Graph Atlas Distance) to the Hallucination Leaderboard [4].
2 The Topologies of LLM Graph Hallucinations
@@
Our aim is to prompt LLMs with famous graphs for which the ground truth topology is known and available online (see e.g. on repositories such as SNAP [14] or KONECT [13]), and consequently most likely part of the training of these LLMs [10]. We chose three graphs in that regard, as ground truth: the Zachary’s karate club graph (coined KC hereafter) and Les Misérables (coined LM, see their composition e.g. in the NetworkX library [8]) and the 50th graph of the graph atlas [18].
Prompts to LLMs
@@
Prompts to LLMs, in order to obtain a graph structure, are in the following form: Provide me the so called "X" graph as a python edge list; print it, with X being the graph of interest, for instance the "Zachary’s karate club", "Les Misérables" or Provide me with graph # from the Graph Atlas, as a python edge list; print it, with # the graph number in case of requests concerning the graph atlas. The request for Python structured responses is because of the NetworkX library [8] we use to instantiate and analyze these graphs.
Outputs by LLMs
@@
Following such prompts, the returned payload is most often in the form of a list such as $(1,2),(1,3),\dots$ (see Section A for an example); that edge list is then parsed and built as a NetworkX graph (undirected ones in alignment with our ground truth graphs in this paper). As these ground truth graphs are also present in the NetworkX library, comparison is convenient. We note that incomplete outputs (i.e. incomplete responses leading to a partial edge list) are nevertheless examined. We report that a small fraction of queried LLMs refused return an edge list, as they claim not having access to the data for instance (as also noted by the Hallucination Leaderboard project [4]); some also prefer to provide Python code to print the queried graph using a library: these models are discarded from our study.
Prompted LLMs
@@
The simplicity and low volume of necessary requests (prompts) enables a full online experience (as opposed to model downloading or API interfacing), using platforms that place LLMs in prompting access via a web browser. We leveraged the following platforms in this paper: Mistral [7], Vercel AI SDK [1], HuggingChat [5], ChatGPT [2], together.ai [9] and Google’s Gemini [3], with the default parameters the platforms set for their hosted models.
Comparing resulting graphs to the ground truth
@@
In this paper we perform topological comparisons, with no consideration for the labels of nodes returned by the LLMs. These are varying significantly; in consequence a hard label matching would discard output graphs that are nevertheless topologically close. We leave a study on the labeling mismatch to future work.
2.1 Statistics on the Topology of Output Graphs by LLMs
@@
Figure 1: Prompting gpt4o for the Zachary’s karate club graph: (a) answered output graph, (b) intersection of the output graph with the KC ground truth graph (a), (c) the edges added (hallucinated) w.r.t. to the KC graph, and (d) the edges missing w.r.t. the KC graph.
We report in detail the prompting of 21 LLMs against the Zachary’s karate club graph (KC); some statistics regarding Les Misérables and graph atlas 50 are deferred in Appendix B. As each prompt to a LLM results in a graph, one can directly perform topological comparison between the ground truth and a LLM output graph; such a comparison is represented in Figure 1: Figure 1a presents the raw output graph; Figure 1b presents the graph intersection of the KC graph with the output graph (from Figure 1a); Figure 1c presents the hallucinated edges (i.e. edges not present in KC but present in the output graph); finally Figure 1d presents the edges that are forgotten in the output graph as compared to KC, while prompting gpt4o. We note that the result provided by this LLM is relatively accurate in comparison to others, as we shall now see in Table 1.
@@
We denote the set $V$ of nodes and $E$ of edges in each examined graph. We name output graphs in relation with their LLM in the first column (in Table 1 the first row being the KC ground truth graph), and provide 6 relevant statistics for assessing their quality [^1] We note that the graph edit distance is not included in these metrics as already intractable for sizes of around 34 nodes as in the output graphs we deal with; this metric is leveraged in Section 3 as we there deal with smaller graphs. We present an alternative and more scalable distance metric in Appendix C.. From left to right, we list the number of nodes $|V|$ in the output graph, its number of edges $|E|$, its density, assortativity (tendency of nodes to attach to nodes with similar degrees), and modularity (that is computed w.r.t. the partition provided by a default label propagation on the output graph [8]). We argue modularity is an important indicator considering the frequent use of KC as a benchmark for community detection methods.
Finally we report in the last column the $l2$ distance between the degree distribution of the output graph and KC. Please note that except for rare unigraphic graphs, there exists multiple graphs having the same degree distribution, so this metric is not a correctness assessment per se. In other words, while a positive distance implies an incorrect result, a distance of 0 does not imply a perfect result (i.e. isomorphic to KC). The table is sorted using this last column.
Table 1 shows that no output graph is exactly correct: all LLMs do hallucinate. The closest one being by the dbrx-instruct model, with just two hallucinated (i.e. added) edges. ChatGPT models (3.5 and 4o) then follow and both return the same output graph. Many outputs are radically different from the correct one (hence deserving being coined hallucinations), with a surprisingly high heterogeneity among the answers (for instance, with $|E|$ ranging from 8 to 153).
@@
The bottom of the list is populated by output graphs that heavily differ from the ground truth, with i.e. the graph on Figure 2a that is a star with two central nodes (resulting in a low assortativity score). In this precise case, the number of nodes is correct, 34, and the number of edges close to correct, showing that model has probably learned these global features but not the topology. Figure 2b demonstrates the opposite case for llama-2-13b-chat-hf, where node and edge counts are very low (6 and 8 respectively). Another interesting case is mistral-large having the correct number of nodes while hallucinating nearly twice the number of edges (see Figure 2c), explaining its high distance to the degree sequence of KC, and its poor modularity in comparison. Finally, Figure 2d showcases for snowflake-arctic-instruct the output of a dense graph (except for a node, resulting in a high density).
Regarding model families, we observe with variants of llama that a higher number of parameters does not imply increased correctness. Indeed, the largest llamma, lama-3.1-405B-Instruct-Turbo, is ranked in the middle. We also observe that for a model version (llama-3-70B), different variants (lite, turbo and groq) result in related ranks, yet with significantly different characteristic outputs.
2.2 Embedding LLMs Output Graph and the Ground Truth
We now rely on graph embeddings, for examining the relative proximity of LLM outputs. We employ the Karate Club library [6] for that purpose, with the NetLSD method [21] with default parameters, before we run a t-SNE appearing in Figure 3. (Les Misérables and graph atlas 50 representations are deferred in Appendix B.)
@@ 2.2 Embedding LLMs Output Graph and the Ground Truth
Despite the proximity of model families such as llama-2 or ChatGPT, a relatively smooth spread appear in this t-SNE, with clusters that mix graphs obtained by different vendors. Older versions of llama models (version 2) appear the less related to KC in this embedding.
3 A LLM Hallucination Rank Based on Output Graphs
Benchmarks for measuring hallucinations are constituted by tens of thousand of prompts [4, 17]. They hence require a privileged access to the model, compared to web browser based prompting access. We here propose to sequentially prompt a LLM for just a handful of graphs from the graph atlas, and to average errors made regarding this reference under a single value (using the average of graph edit distances).
The graph atlas is composed by 1252 different graphs; in our experiment, we choose a resolution of $5$ graphs to be prompted, in particular the first $5$ connected graphs (namely graphs #3, #6, #7, #13 and #15). We then compute the exact edit distance of each of these ground truth graphs against their respective LLM output. We finally average these $5$ distances to obtain the final distance score for the queried LLM. We coin this distance the Graph Atlas Distance (GAD).
We note that across all tested LLMs and for all datasets, an isomorphism solely occurred for gpt4o on graph atlas #7 (a triangle graph) and #13 (a star composed of 4 nodes). The weight given to edit operations are is follows: we do not account for labels, but consider node/edge insertion/deletion each costing 1.
@@ 3 A LLM Hallucination Rank Based on Output Graphs
Figure 4: GAD (resolution: 5 graphs) vs a summarization of a benchmark based on prompts with binary answers [4]. In the GAD column, the first number is the average edit distance distance, the second is the standard deviation. ($\star$ rank in [4] as been adjusted to the LLMs available in our experiment).
@@
We can observe an interesting correlation in these two rankings (with a spearman rank correlation of $0.3$, where 00 is random). The first position is held by gpt4o. The first of the 4 llama models (llama-3.1-70B-Instruct) is down-ranked in GAD, yet the 3 others are in the same order. The larger 405B parameters llama perform worse than smaller models; note that this inversion also appears in [4], where llama 3 beats llama 3.1 with the same amount of parameters (70B). qwen2-72B-Instruct similarly ranks in the middle. c4ai-command-r-plus ranks at the bottom. snowflake-arctic-instruct is nevertheless strongly down-ranked with GAD, as compared to its second position in the Hallucination Leaderboard.

### v3

# LLMs Prompted for Graphs: Hallucinations and Generative Capabilities

@@ Abstract
Large Language Models (LLMs) are nowadays prompted for a wide variety of tasks. In this article, we investigate their ability in reciting and generating graphs. We first study the ability of LLMs to regurgitate well known graphs from the literature (e.g. Karate club or the graph atlas) [^1] This article is an extended version of the conference paper [22].. Secondly, we question the generative capabilities of LLMs by asking for Erdős–Rényi random graphs. As opposed to the possibility that they could memorize some Erdős–Rényi graphs included in their scraped training set, this second investigation aims at studying a possible emergent property of LLMs. For both tasks, we propose a metric to assess their errors with the lens of hallucination (i.e. incorrect information returned as facts). We most notably find that the amplitude of graph hallucinations can characterize the superiority of some LLMs. Indeed, for the recitation task, we observe that graph hallucinations correlate with the Hallucination Leaderboard, a hallucination rank that leverages $10,000$ times more prompts to obtain its ranking. For the generation task, we find surprisingly good and reproducible results in most of LLMs. We believe this to constitute a starting point for more in-depth studies of this emergent capability and a challenging benchmark for their improvements. Altogether, these two aspects of LLMs capabilities bridge a gap between the network science and machine learning communities.
@@
Large language models, graphs, hallucinations, Erdős–Rényi, random generation.
@@ 1 Introduction
Large Language Models (LLMs) recently attracted a lot of attention, thanks to sustained research efforts and a large spectrum of envisioned applications. This triggers a growing demand for tools to test and analyze these complex and expensive-to-set-up objects. In particular, methods to efficiently identify, differentiate, watermark LLMs, and methods to compare their accuracy and notably the potential presence of hallucinations are devised [25, 32, 30]. The common denominator of these methods is the will to efficiently extract information from the LLM under scrutiny.
@@
To achieve this information extraction, one can distinguish two broad categories of approaches: white box approaches that rely on exploiting a privileged access to the model internals (e.g. probits, activation patterns, source code, model hyperparameters and weights), and black box approaches that merely allow an auditor to interact with a given model. White box approaches provide precise answers, but are not always feasible, for instance, to an external auditor assessing a closed source LLM. In such cases, black box approaches are necessary and constitute the focus of this article.
@@
To conduct a black box audit of a LLM, a common approach is to rely on multiple choice questions (MCQ) and test the accuracy of its answers. Using MCQ circumvents the difficulty and subjectivity of evaluating (a potentially large number of) answers to open questions. To this end, several MCQ datasets are readily available, and cover a wide range of knowledge and reasoning domains. For instance, MedMCQA [28] contains 182k four-choice questions assessing the target model general medical knowledge. Using this dataset yields 2 bits of information on the target model per requested prompt. This might reveal inefficient if the auditor has to pay for each prompt or if she has to limit the rate at which prompts are sent.
@@
In this article, we explore a different approach using prompts that aims at obtaining graphs rather than text. As graphs are structured, requesting graphs circumvents the problem of analyzing free form text answers. And most importantly, a single request to obtain a $n$ node graph from a LLM yields $\mathcal{O}(n^{2})$ structured information bits from that model, enabling a much more efficient information extraction rate per request for the auditor. For instance, requesting the standard “Les Misérables" graph is much more efficient than issuing queries like "do Jean Valljean and Cosette both appear in a chapter?", for any two of the 77 characters of the book. Our first research question, RQ1, focuses on studying the LLM ability to recite graphs they most likely have viewed during their training. Indeed, throughout the years, network science has produced a number of standard benchmark graphs that are probably already ingested by most LLMs and that provide a readily available ground truth, e.g. in the SNAP [23] or KONECT [21] repositories among others), and as the training data of LLMs is known to include most of online scrapable data [13]. While with RQ1 we question to which extent LLMs are able to memorize given examples of graphs, with our second research question RQ2, we interrogate a very different ability, generating randomness in a structured form. We choose Erdős–Rényi random graphs. RQ2 is inspired by studies on randomness generations with LLMs [17, 20, 16] and on studies on LLMs emergent abilities. In the literature, an emergent ability can be defined as an ability to reason about novel problems in a zero-shot setting, i.e. without any training on the sought task [35].
Though querying LLMs for graph structures is unexplored and might sound promising, it raises a number of questions we further investigate in the following. First, we detail the interaction framework in which the objective is to collect a graph as a LLM output. Once this has been established, we focus on the quality of the collected graphs proposing adapted metrics to answer RQ1 and RQ2, within the lens of hallucination. Hallucination is commonly defined as when the generated content is nonsensical or unfaithful to the provided source content [19, 39]. In RQ1, as there is a ground truth comparison, any error can be interpreted as hallucination. In RQ2, we measure the frequency with which LLMs produce highly inaccurate random graphs, referring to such errors as hallucinations.
Finally, we showcase the potential of graph queries and their ability to extract information from LLMs in a request-efficient manner. To this end, we compare the Hallucination Leaderboard [1] rank to the rank we produce with our two proposed metrics.
2 The Topologies of LLM Graph Hallucinations (RQ1)
@@
Our aim is to ask LLMs for famous graphs for which the ground truth topology is commonly available online (see e.g. on repositories such as SNAP [23] or KONECT [21]), and consequently most likely part of the training set used for these LLMs [13]. We chose three graphs, the Zachary’s Karate Club graph (coined KC hereafter) and Les Misérables (coined LM, see their description e.g. in the NetworkX library [8]) and the 50th graph of the graph atlas [29].
Prompting LLMs.
@@
The prompts we used in order to obtain a graph structure are the following: Provide me the so called ‘‘X" graph as a python edge list; print it, with X being the graph of interest, for instance the KC or the LM. We also used the prompt Provide me with graph # from the Graph Atlas, as a python edge list; print it, with # being the graph number in the case of queries concerning the graph atlas. We asked for Python structured responses because we use the NetworkX library [8] to instantiate and analyze the outputs.
LLM Outputs.
@@
Given the chosen prompts, the outputs are most often in the form of a list, e.g. $(1,2),(1,3),\dots$ (see Section A for an example). The edge list is then parsed and built as a NetworkX graph; we used undirected graphs in alignment with the ground truth graphs chosen in this article. As the ground truth graphs are also present in the NetworkX library, the comparison is convenient. To be noted that incomplete outputs (i.e. incomplete responses leading to partial edge lists) are nevertheless examined. We report that a small fraction of queried LLMs refused to return an edge list and claimed they do not have access to the data (as also noted by the Hallucination Leaderboard project [1]). Some other LLMs only provide the Python code to print the queried graph using a certain library (e.g. NetworkX); these models are discarded from our study.
Prompted LLMs.
@@
The simplicity and low volume of necessary queries (prompts) enables a non-automated and full online experience (as opposed to downloading the model or using the API interface). We used online platforms that allow prompting access to LLMs via a web browser. In this work, we used the following platforms: Mistral [7], Vercel AI SDK [10], HuggingChat [5], ChatGPT [3], together.ai [9] and Google’s Gemini [4], with the default parameters the platforms defined for their hosted models.
Comparing output graphs to the ground truth.
@@
In RQ1 we perform topological comparisons without considering the node labels returned by the LLM (e.g. characters names for the Les Misérables graph). These vary significantly, and thus a hard label matching would discard output graphs that might be topologically close. We leave a study on the labeling mismatch to future work.
2.1 Statistics on the Topology of Output Graphs
@@
Figure 1: Prompting gpt4o for the Zachary’s karate club graph. (a) answered output graph, (b) intersection of the output graph with the KC ground truth graph, (c) the edges added (hallucinated) w.r.t. to the KC graph, and (d) the edges missing w.r.t. the KC graph.
We report in detail the prompting of 21 LLMs when we asked for a Zachary’s Karate Club (KC) graph. Some other statistics on Les Misérables and graph atlas 50 are deferred in Appendix B. As each prompt results in a graph, one can directly perform topological comparison between the ground truth and a LLM output graph. Such a comparison is represented in Figure 1. Figure 1a shows the raw output graph, while Figure 1b shows the graph intersection of the KC graph with the output graph from Figure 1a. Figure 1c presents the hallucinated edges (i.e. edges not present in KC but present in the output graph). Finally, Figure 1d presents the edges that are forgotten in the output graph compared to the KC, when prompting gpt4o. We note that the result provided by gpt4o is relatively accurate in comparison to the others, as we can also see in Table 1.
@@
For each examined graph, we denote the set of nodes and of edges with $V$ and $E$ respectively. We name output graphs in relation with their generating LLM in the first column (in Table 1 the first row being the KC ground truth graph), and provide six relevant statistics for assessing their quality [^2] We note that the graph edit distance [14] is not included in these metrics as it becomes intractable for sizes of around 34 nodes, as it can be seen in the output graphs of our sample. This metric is leveraged in Section 3 as we deal with smaller graphs. We present an alternative and more scalable distance metric in Appendix C.. From left to right, we list the number of nodes $|V|$ in the output graph, its number of edges $|E|$, its density, assortativity (i.e. tendency of nodes to attach to nodes with similar degrees), and modularity (that is computed w.r.t. the partition provided by a default label propagation on the output graph [8]). We argue that modularity is an important indicator considering the frequent use of KC as a benchmark for community detection methods.
Finally we report in the last column the $L2$ distance between the degree distribution of the output graph and KC. Please note that except for rare unigraphic graphs, there exists multiple graphs having the same degree distribution, so this metric is not a correctness assessment per se. In other words, while a positive distance implies an incorrect result, a distance of 0 does not imply a perfect result (i.e. isomorphic to KC). The table is sorted using the last column.
Table 1 shows that no output graph is exactly correct, i.e. all LLMs hallucinate. The closest being the dbrx-instruct model, with just two hallucinated (i.e. added) edges. ChatGPT models (3.5 and 4o) follow and both return the same output graph. Many outputs are radically different from the correct one, with a surprisingly high heterogeneity among the answers (for instance, with $|E|$ ranging from 8 to 153).
@@
The bottom of the list is populated by output graphs that heavily differ from the ground truth, e.g. the graph on Figure 2a that is star shaped with two central nodes (resulting in a low assortativity score). In this precise case, the number of nodes is correct, 34, and the number of edges close to correct, showing that model has probably learned these global features but not the topology. Figure 2b demonstrates the opposite case for llama-2-13b-chat-hf, where node and edge counts are very low (6 and 8 respectively). Another interesting case is mistral-large, which generates the correct number of nodes while hallucinating nearly twice the number of edges (see Figure 2c), which explains its large distance to the degree sequence of KC, and its poor modularity in comparison. Finally, Figure 2d showcases for snowflake-arctic-instruct the output of a dense graph (except for a node, resulting in a high density).
When considering families of models, we observe that in the case of the various versions of Llama, a higher number of model parameters does not imply increased correctness. Indeed, the largest Llama, lama-3.1-405B-Instruct-Turbo, has a middle rank. We also observe that for a model version (llama-3-70B), different versions (lite, turbo and groq) result in related ranks, yet with significantly different characteristic outputs.
2.2 Embedding LLMs Output Graph and their Ground Truth
We now rely on graph embeddings to evaluate the relative proximity of LLM outputs. We employ the Karate Club library [6] with the NetLSD method [33] with default parameters and Figure 3 shows their t-SNE (Les Misérables and graph atlas 50 representations are deferred in Appendix B.)
@@ 2.2 Embedding LLMs Output Graph and their Ground Truth
Despite the proximity of model families such as Llama-2 or ChatGPT, a relatively smooth spread appears in the t-SNE, with clusters that mix graphs belonging to different vendors. Older versions of Llama models (version 2) appear the less related to KC in this embedding.
3 A LLM Hallucination Ranking Based on Output Graphs
In this section, we continue to refer to RQ1 and we propose a dedicated metric which allows to rank LLMs based on their hallucination amplitude. On the other hand, benchmarks for measuring hallucinations are constituted by tens of thousand of prompts [1, 28] and they hence require a privileged access to the model, compared to web browser based prompting access. Here we propose to sequentially prompt a LLM for just a handful of graphs from the graph atlas, and to average errors made with respect to the reference under a single value (that is the average of graph edit distances).
The graph atlas is composed by $1,252$ different graphs. In our experiment, we choose a resolution of $5$ graphs to be prompted, in particular the first $5$ connected graphs (namely graphs #3, #6, #7, #13 and #15). We then compute the exact edit distance of each of these ground truth graphs against their respective LLM output. We finally average these $5$ distances to obtain the final distance score for the queried LLM. We refer to this distance as the Graph Atlas Distance (GAD).
We note that across all tested LLMs and for all datasets, an isomorphism solely occurred for gpt4o on graph atlas #7 (a triangle graph) and #13 (a star shaped graph composed of 4 nodes). The weight given to edit operations are as follows: we do not account for labels, but consider node/edge insertion/deletion, each costing 1.
@@ 3 A LLM Hallucination Ranking Based on Output Graphs
Table 2: GAD (resolution: 5 graphs) vs a summarizing of a benchmark based on prompts with binary answers [1] (ranking from 2024-08-20). In the GAD column, the first value shows the average edit distance distance, the second the standard deviation. ($\star$ rank in [1] as been adjusted to the LLMs available in our experiment).
@@
We can observe an interesting yet rather weak correlation in these two rankings (with a Spearman rank correlation of $0.3$, where 00 is random). The first position is held by gpt4o. The first of the four Llama models (llama-3.1-70B-Instruct) is down-ranked in GAD, yet the three others are in the same order. The larger 405B parameters Llama perform worse than smaller models (to be noted that this inversion also appears in [1], where Llama 3 beats Llama 3.1 with the same amount of parameters (70B)). qwen2-72B-Instruct similarly ranks in the middle, while c4ai-command-r-plus ranks at the bottom. snowflake-arctic-instruct is nevertheless strongly down-ranked with GAD, as compared to its second position in the Hallucination Leaderboard.